One gram of sodium hydroxide pellets is dissolved in two hundred-fifty milliliters water. What is the molarity, molality and normality of the solution? How much final solution is formed in terms of grams, milliliters and total number of moles? Assume the density of solution to be approximately that of water and the density of sodium hydroxide pellets to be 2.13 g/cc.( 4 decimal places)


1
Expert's answer
2021-06-01T02:44:39-0400

Number of moles =1÷40=0.025 moles


Molarity ="\frac {0.025}{0.25}=0.1M"


Final solution=24g,267.6mL